OpenStudy (anonymous):

Prove or disprove the following statements. I and 0 denote respectively the identity and zero matrix of the same size as A. (a) A system with three equations and two unknowns must be inconsistent. (b) A system of n equations with n unknowns has at most n solutions. (c) AandBare2×2matricessuchthatAB=0thenwemusthaveA=0orB=0 (d) If the system AX = 0 has infinitely many solutions, so does the system AX = b for any choice of b. (e) If A is a square matrix such that A2 −3A+2I = 0 then A−cI is invertible whenever c ̸= 1 and c ̸= 2. Please help, I am very lost!

OpenStudy (loser66):

@e.mccormick you should be here

OpenStudy (e.mccormick):

Well, one way to disprove things is to show an example that contradicts the statement.

OpenStudy (loser66):

those information is basic, right? need to understand? make question and get explanation, please

OpenStudy (e.mccormick):

Is this inconsistent or consistent: \(2x +y =4\) \(x +y =3\) \(2x +2y =6\)

OpenStudy (anonymous):

Thanks I just got it!

OpenStudy (e.mccormick):

A lot of those are just applying the rules. For each rule there is a proof. So you just need to find each proof.